1) Find the impulse response of de VCE) + 6 mV (t) + 5y(t) = x(t) + 3x(t)
Answer #1

ce + 6 (+57(4) = 4*4 + 3x(4) Find Impulse Response of the system > Impulse Response of a system is response when input is 8(tY(s) = _5+3 (5+5)(st) :: $2+65+5 = (3+5) (sti) Apply Partial fraction is+5)(s+I) (s+5) IST → (3+3) = A (5+1)+ B (+) - At s= -so y(s) = 1/2 (sts) + 2 (st) sta Now I[ k] = ke duley Alpbly Laploce Invese to Y15) YC0 = ý cet & Jure J (0 = 0-5 [€ 54 e tj